OpenStudy (anonymous):

Solve. 3^x-5 = 9 also how do i find what is x?

Parth (parthkohli):

Equate the bases please. \(\Large \color{MidnightBlue}{\Rightarrow 3^{x - 5} = 3^2 }\) Equate the exponents. \(\Large \color{MidnightBlue}{\Rightarrow x - 5 = 2 }\) \(\Large \color{MidnightBlue}{\Rightarrow x = 7 }\)

OpenStudy (anonymous):

why 2?

Parth (parthkohli):

I have made an equation of the exponents.

OpenStudy (anonymous):

ohh

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow a^m = a^n \Rightarrow m = n }\)

OpenStudy (anonymous):

7^x - 5 = 74^x + 7 so what about that one?

OpenStudy (anonymous):

Wait 9+5=14 so wouldn't whatever x be have to end up at 14?